OpenStudy (anonymous):

Verify each identity: (sin^2x-cos^2x)/(sinx+cosx)=sinx-cosx

OpenStudy (anonymous):

\[\frac{ \sin ^{2}x-\cos ^{2}x }{ \sin x+\cos x }=\frac{ \left( \sin x+\cos x \right)\left( \sin x-\cos x \right) }{ \sin x+\cos x} =?\]

OpenStudy (anonymous):

ohhh so we just cross it out?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks!!!!!!!! :)

OpenStudy (anonymous):

@alessandro96